OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

OpenStudy (anonymous):

x=-6x+18=-12

OpenStudy (anonymous):

-12-18=-22

OpenStudy (anonymous):

is that the final answer?

OpenStudy (anonymous):

? oh, soory, -30./6=-5

OpenStudy (anonymous):

the final answer is 5 not -5

OpenStudy (anonymous):

u put 6 instead of -6

OpenStudy (anonymous):

ok. do you have tha answer now?

OpenStudy (anonymous):

it is -5. but alright. were done here

OpenStudy (anonymous):

the absolute sign prohibits negetive 6, but do as you will.

OpenStudy (tkhunny):

−3|2(-5) + 6| = -3|-10 + 6| = -3|-4| = -3(4) = −12 Good. x = -5 checks out. −3|2(-1) + 6| = -3|-2 + 6| = -3|4| = -3(4) = −12 Whoops. Looks like you'all missed one. x = -1. How could you have avoided that?

OpenStudy (tkhunny):

@hermineb @Fireflame13

OpenStudy (anonymous):

thxx Tkhunny

OpenStudy (anonymous):

thxx fireflame13 but u wasnt specific thats y i was confused

OpenStudy (anonymous):

it's fine.

OpenStudy (tkhunny):

Let's try it again: -3|2x + 6| = −12 Stuff outside the the Absolute Values doesn't do much. Let's make our lives simpler. |2x + 6| = 4 Common factor in the two terms inside the Absolute Values. Let's make our lives simpler. |2x + 6| = |2(x+3)| = 2|x+3| = 4 and |x+3| = 2 That's WAY easier. x+3 = 2 or x+3 = -2 And both solutions are easily found.